OpenStudy (anonymous):

A refrigerator has a volume given by the trinomial y3 + y2 – 56y. What are the possible dimensions of the refrigerator? Use factoring. A. y, y + 8, and y – 7 B. y, y – 8, and y + 7 C. y + 7, y + 7, and y – 7 D. y + 8, –y + 8, and –y + 7

OpenStudy (zehanz):

Hey @XbriX again! This one looks much like the other. Only now you can first factor y: y(y²+y-56) = y(y+a)(y+b) Now remember what you learned earlier: you are looking for a and b, with a+b=1 and a*b=-56 What are a and b?

OpenStudy (anonymous):

y, y – 8, and y + 7?(:

OpenStudy (zehanz):

-8+7=-1 :( -8*7=-56 :) So it is 8 and -7...

OpenStudy (anonymous):

y, y + 8, and y – 7?

OpenStudy (zehanz):

Right!

OpenStudy (anonymous):

would u help me with another? A sphere has a radius of 4x + 4. Which polynomial in standard form best describes the total surface area of the sphere? Use the formula S = 4pir2 for the surface area of a sphere. A. 64pix2 + 128pix + 64pi B. 64pix2 – 128pix + 64pi C. 64pix2 + 128pix – 64pi D. 64pix2 – 128pix – 64pi

OpenStudy (zehanz):

Put the radius in the formula: 4π(4x+4)² and expand...

OpenStudy (zehanz):

(first, do (4x+4)², then multiply everything with 4π.

OpenStudy (anonymous):

64pix2 + 128pix + 64pi?

OpenStudy (zehanz):

Right again!

OpenStudy (anonymous):

are you serious? im actually starting to understand it lol. thank you!